### v1

# The distribution of misalignment angles in multipolar planetary nebulae

Abstract
We measure the projected angle on the plane of the sky between adjacent symmetry axes of tens of multipolar planetary nebulae and find that the distribution of these misalignment angles implies a random three-dimensional angle distribution limited to $\lesssim 60^{\circ}$. We identify a symmetry axis as a line connecting two opposite lobes (bubbles) or clumps. We build a cumulative distribution function of the projected angles $\alpha$ and find that an entirely random distribution of the three-dimensional angles $\delta$ between adjacent symmetry axes, namely, uncorrelated directions, does not fit the observed one. A good fit to the observed distribution is a limited random distribution of the three-dimensional angle between adjacent symmetry axes, i.e., random distribution in the range of $20^{\circ}\lesssim\delta\lesssim 60^{\circ}$. We assume that a pair of jets along the angular momentum axis of an accretion disk around the companion shape each symmetry axis. The limited random distribution might result from two sources of angular momentum to the accretion disks with comparable magnitude: one with a fixed direction and one with a stochastic direction variation. We discuss a scenario where the fixed-axis angular momentum source is the binary orbital angular momentum, while the stochastic source of angular momentum is due to the vigorous envelope convection of the mass-losing giant progenitor.
stars: jets – stars: AGB and post-AGB – binaries: close – stars: winds, outflows – planetary nebulae: general
1 Introduction
Most planetary nebulae (PNe) have one or more pairs of structural features on opposite sides of the center. The structural features include bubbles (a bubble is a faint zone closed by a bright rim), lobes (a lobe is a faint zone with a partial outer bright rim open to the far side), clumps (also called ansae), nozzles (a nozzle is a narrow opening in the PN shell), ears (an ear is a protrusion with a base smaller than the PN main shell and a cross-section that decreases outward), or arcs (rims). Hereafter, we refer to the two opposite structural features as poles and the line connecting them as a symmetry axis. PNe that have two or more symmetry axes are termed multipolar PNe (e.g., Sahai & Trauger 1998; Steffen et al. 2013; Huang et al. 2016; Rechy-García et al. 2020; Bandyopadhyay et al. 2023; Wen et al. 2023. Multipolar PNe with two symmetry axes are called quadrupole PNe (e.g., Manchado et al. 1996b; Guerrero et al. 2013).
Multipolar PNe have been the center of some past studies (see the recent review by Kwok 2024), e.g., Velázquez et al. (2012) modeled precessing jets with varying velocity as the cause of multipolar PNe. Chong et al. (2012) suggested that multipolar structures are common in PNe, much more common than what is directly observed (without analysis). In this study, we will study only the images of multipolar PNe, referring only to their appearance on the plane of the sky.
Most studies attribute each symmetry axis to a pair of jets launched by a binary system progenitor (e.g., Morris 1987; Soker 1990; Sahai & Trauger 1998; Akashi & Soker 2018; Estrella-Trujillo et al. 2019; Tafoya et al. 2019; Balick et al. 2020; Rechy-García et al. 2020; Clairmont et al. 2022; Danehkar 2022; Moraga Baez et al. 2023; Derlopa et al. 2024; Miranda et al. 2024; Sahai et al. 2024 for a list of a small fraction of all papers; Baan et al. 2021 present an alternative scenario based on fallback material). In most cases, the binary system progenitor experiences a common envelope evolution of a main sequence star orbiting inside the envelope of an asymptotic giant branch (AGB) star; a red giant branch (RGB) star as the primary star is also possible (e.g., Hillwig et al. 2017; Sahai et al. 2017; Jones et al. 2020, 2022, 2023).
The pair of jets might change direction during the PN formation phase. One pattern is the precession of the jet axis, as observed in some PNe (e.g., Guerrero & Manchado 1998; Miranda et al. 1998; Sahai et al. 2005a; Boffin et al. 2012; Sowicka et al. 2017; Rechy-García et al. 2019; Guerrero et al. 2021; Clairmont et al. 2022). Another pattern is jet-launching episodes with non-monotonic varying axis directions. Soker (2023a) suggests that most of the brightest PNe are multipolar formed in a violent process with non-continuous direction variation of the jets’ axis. We return to this point in our discussion in Section 4.
In this study, we examine multipolar PNe, where different pairs of lobes indicate separate jet-launching episodes rather than precession or another continuous change of direction. In Section 2, we present the sample of multipolar PNe with the projected angles between their symmetry axes. In Section 3 we analyze the angle distribution and summarize our results in Section 4
2 The sample of multipolar PNe
In Tables 1 and 2, we list the 40 PNe and 5 pre-PNe, respectively, for which we could find two or more well-defined symmetry axes. In the second column, we give the angle between the axes in each PN (in degrees), referencing an image in the last column.
Table 1: Planetary nebulae with binaries and jets
PN $\alpha$ Ref
M 1-16 6 GM23
M 2-46 8 Ma96
Hen 2-96 8 We24
M 1-61 12; 55 Sa11
J320 15; 18 Ha04
Pe 1-1 15; 23; 48 We23
Hen 2-73 16 We24
Hen 2-158 17; 47 We23
NGC 6790 18 Hs14
NGC 6886 18 PNIC
NGC 6072 19; 38 Kw10
NGC 6644 20 Sc92
He 2-47 21; 22; 46 Sa00
Hen 2-447 24 PNIC
M 1-75 25 Ma96
M1-37 27; 36 Sa00
Me 2-2 27; 48 Sa98
IC 5117 29 Hs14
NGC 6309 30 Ru15
NGC 5307 31; 46 PNIC
PN $\alpha$ Ref
NGC 6302 32 Sc92
IC 4634 33 Gu08
NGC 6572 33 Ak16
M 1-31 40; 42 PNIC
Kn 26 41 Gu13
IC 4846 42 Mi01
NGC 7026 46 Cl13
KjPn 8 49 Lo00
M 4-14 51 Ma96
M 1-33 52 PNIC
NGC 6445 56 Sc92
NGC 2440 58 Lo98
M 3-28 59 Ma96
NGC 7027 61 Mo23
Hen 2-115 61 Sa98
NGC 2371 63 GG20
M 3-35 72 PNIC
K 3-24 74 Ma96
NGC 5315 80 PNIC
M 1-59 82 Go24
Note: Projected angle between jet axes in multipolar PNe (second column in degrees).
Abbreviation: PN: planetary nebula; Ref: image reference.
References: Ak16: Akras & Gonçalves (2016) and Bandyopadhyay et al. (2023); Cl13: Clark et al. (2013); GG20: Gómez-González et al. (2020); GM23: Gómez-Muñoz et al. (2023); Go24: Gold et al. (2024); Gu08: Guerrero et al. (2008); Gu13: Guerrero et al. (2013); Ha04: Harman et al. (2004); Hs14: Hsia et al. (2014); Kw10: Kwok et al. (2010); Lo98: López et al. (1998); Lo00: López et al. (2000); Ma96: Manchado et al. (1996a); Mi01: Miranda et al. (2001); Mo23: Moraga Baez et al. (2023); PNIC: HST archive images from the Planetary Nebula Image Catalogue (PNIC) of Bruce Balick (Balick, 2006); Ru14: Rubio et al. (2015); Sa98: Sahai & Trauger (1998); Sa00: Sahai (2000); Sa11: Sahai et al. (2011); Sc92: Schwarz et al. (1992); We23: Wen et al. (2023); We24: Wen et al. (2024).
Table 2: Misalignment angles in multipolar pre-PNe
PN $\alpha$ Ref
IRAS04395+3601 15 Tr02
IRAS16594-4656 17; 23 Hr99
IRAS19024+0044 32 Sa05
IRAS17047-5650 47 Sa07a
IRAS19475+3119 62 Sa07b
Note: Similar to Table 1 but for proto-PNE.
References: Hr99: Hrivnak et al. (1999); Sa05: Sahai et al. (2005b); Sa07a: Sahai et al. (2007a); Sa07b: Sahai et al. (2007b); Tr02: Trammell & Goodrich (2002);
We present several specific examples. The first is PN M 2-46. We present its image in Figure 1, adapted from Manchado et al. (1996a), as an example of a case with one pair of lobes entirely inside the other pair.
Figure 1: An image of the multipolar (Quadrupolar) PN M 2-46 adapted from Manchado et al. (1996a) with our marks in red. This is an example of one pair of lobes entirely inside the other.
In Figure 2 we present an image of the PN KjPn 8 adapted from López et al. (2000). They already marked the two symmetry axes (black lines); we mark the projected angle between the two axes $\alpha$. In the PN KjPn, the small pairs of lobes (A1-A2) are almost entirely inside the large pairs of lobes (C1-C2). In addition, we mark three rims in the eastern lobe with red arrows. Three jets probably compressed these three rims; we do not find evidence of three jet-launching episodes in the western lobe. These three rims without counterparts in the opposite lobe suggest that our analysis might miss some jet-launching episodes with a slight misalignment angle. We return to this point in Section 3.
Figure 2: An image of the multipolar PN KjPn 8 adapted from López et al. (2000) with their marks in black and our marks in red. The three rims suggest three jet-launching episodes that inflated the large lobes. In this study, we consider them to be one event.
Figure 3 presents the PN J320, where we identify three symmetry axes, two with pairs of clumps and one with a pair of bubbles. We present this PN to demonstrate a case where the location of the poles is not well defined and where the symmetry axes that we draw by the poles miss the center somewhat. The uncertainties in the projected angles are about $2^{\circ}$. Because we study the cumulative distribution function (Section 3), such uncertainties have negligible influence on our conclusions.
Figure 3: An image of the multipolar J320 adapted from Harman et al. (2004) with our marks in red. This PN demonstrates uncertainties in the exact poles: the location of the four clumps forming two pairs and the two bubbles of the third pair. The lines we chose by the bright zone of the poles of each axis do not cross exactly at the center. Angles between axes are in degrees. The uncertainties here are about $2^{\circ}$. However, these do not change our conclusions on the best statistical angle distribution.
The PN M1-37 that we present in Figure 4 demonstrates a case with two lobes very close to each other, the two in the south. Since there is only one lobe on the other side, the north, we take the two south lobes to belong to the same jet-launching episode. However, we may miss a fourth pair very close to the north-south line that Sahai (2000) drew. We draw a line from the edge of the north lobe through the center; the other side of the line falls between the two neighboring south lobes; for the two other symmetric axes, we take Sahai (2000) marked. This further suggests we might miss small projected angles, $\alpha\lesssim 15^{\circ}$.
Figure 4: An image of the multipolar M1-37 adapted from Sahai (2000). The original marks are in black, and ours are in pale blue. Sahai (2000) marks three symmetry axes. He takes the large lobe in the south. We consider the two lobes in the south to be related to the same jets-launching episode and draw a line from the north lobe through the center of the PN. We agree on the other two axes. This again demonstrates 1-2 degrees of uncertainties in the angles.
Some PNe that studies (e.g., Guerrero et al. 2020) find multipolar do not have clear morphologies for us to measure the angles from the images alone, e.g., IC 4776 (Rechy-García et al., 2020), or are too messy, e.g., NGC 5189 (Sabin et al., 2012) and M 1-26 that might even have three axes, but only one is well defined (image in, e.g., Rechy-García et al. 2020). NGC 6210, which is in the list of Guerrero et al. (2020) of multipolar PNe, is a messy PN that has a precessing jet pair (e.g., Rechy-García et al. (2020) that we do not consider in this study. Guillén et al. (2013) study the multipolar PN NGC 6058. We do not include this PN because we could not identify clear symmetry axes projected on the plane of the sky. Future studies should add more multipolar to our study, those classified by three-dimensional analysis rather than the image on the plane of the sky.
3 The distribution of misalignment angles
We build the cumulative distribution function $W_{\alpha}(\alpha)$, which is the fraction of objects with a projected angle $\alpha_{i}$ in the range of $0\leq\alpha_{i}\leq\alpha$. In Figure 5 we present this function by the thick-black line (the step function). With a green-dotted line, we show a theoretical $W_{\alpha}$ for the case where the two jet axes of adjacent axes are entirely random, i.e., uncorrelated. A fully random distribution does not fit the observed distribution. We have tried a constant three-dimensional angle between any two axes. The best fit is for $\delta_{f}=43^{\circ}$, as shown by the solid-orange line in Figure 5. We compare two other theoretical cases with $\delta=\delta_{f}-5^{\circ}=38^{\circ}$ and $\delta=\delta_{f}+5^{\circ}=48^{\circ}$.
Figure 5: Cumulative distribution functions $W_{\alpha}$ of projected angels between symmetry axes. The black step function is $W_{\alpha}$ from the data of 40 PNe and 5 pre-PNe with 58 angles. The green-doted line is $W_{\alpha}$ for a fully random distribution of angles in their dimension $\delta$; namely, the two or three symmetry axes do not correlate. The three other lines are $W_{\alpha}$ for a constant three-dimensional angle between the two axes, with values given in the inset. The $\delta=43^{\circ}$ is the best fit for a constant value of $\delta$.
We measure the maximum vertical distance $D_{\rm max}$ of each theoretical line from the data. The best fit theoretical line is the one with the minimum value of $D_{\rm max}$. Because we might miss very small projected angles, as the two jet-inflated lobes/ears might merge, we do not include angles of $\alpha\lesssim 15^{\circ}$ in measuring $D_{\rm max}$. The best fit for a constant value of three-dimensional angle $\delta$, i.e., $\delta=43^{\circ}$, has $D_{\rm max}=0.115$ at $\alpha=33^{\circ}$.
We next build a theoretical function with a random distribution of three-dimensional angles $\delta$ within a range bounded by a minimum allowed angle $\delta_{d}$ and a maximum allowed angle $\delta_{u}$. We find the best fit to the observed distribution to be $\delta_{d}=22^{\circ}$, and $\delta_{u}=60^{\circ}$, i.e., $\delta$ is entirely random in the range of $22^{\circ}\lesssim\delta\lesssim 60^{\circ}$; this is the $W_{\alpha}(22,60)$ function. Figure 6 presents this case with a solid-orange line, which has $D_{\rm max}=0.056$ at $\alpha=63^{\circ}$. For comparison, we present four other cases in Figure 6, having larger values of $D_{\rm max}$. The bounded random distribution fits very well with the observed distribution, much better than a constant $\delta$ distribution that we present in Figure 5. Again, we might miss small angles in observed PNe, and we do not include the range of $\alpha\lesssim 15^{\circ}$ in determining the best fit.
Figure 6: Cumulative distribution functions $W_{\alpha}$ of projected angels between symmetry axes. The black step function and the green-doted lines are the data and fully random distribution, respectively, as in Figure 5. The other five lines are for a random distribution of the three-dimensional angle $\delta$ limited by lower and upper bounds, as the inset indicates. The solid-orange line depicts the best fit.
One possible explanation for the very good match of the limited random function $W_{\rm\alpha}(22,60)$ with the data and the mismatch of the fully random function (green dotted line) with the data is the presence of two sources of angular momentum with comparable magnitude. An accretion disk launches two opposite jets along its angular momentum axis. Consider one source of angular momentum with a constant direction and one completely or almost completely random source. If they are of comparable magnitude, or the random is somewhat smaller than the fixed angular momentum component, the two angular momenta add to a new angular momentum direction within a few tens of degrees from the fixed angular momentum axis. The angles between the axes of two consecutive jet-launching episodes will have a random component within a range smaller than $90^{\circ}$.
4 Summary and discussion
We examined images of tens of PNe and pre-PNe. We measured the projected angle on the plane of the sky, $\alpha$, between adjacent symmetry axes that researchers attribute to shaping by pairs of opposite jets (Tables 1 and 2). By the inspection of images without deeper analysis, like velocity measurement, we might miss cases with small projected angles $\alpha$, as some PNe suggest (Figures 2 and 4). This study does not refer to precessing jets continuously changing direction; we only consider separated symmetry axes.
We draw the observed cumulative distribution function $W_{\alpha}$ in Figures 5 and 6 (black step function). We find that an entirely random distribution of the angle $\delta$ between symmetry axes in three dimensions (green-dotted line) does not fit the observed distribution. We plot two other theoretical cumulative distribution functions, one of a constant three-dimensional angle $\delta$ between the two symmetry axes (Figure 5), and one of a random distribution between two angles (Figure 6). Not considering small angles of $\alpha\lesssim 15^{\circ}$ that we might miss in observations, we find the best fit among these functions to be $W_{\alpha}(22,60)$, a limited random distribution between $22^{\circ}$ and $60^{\circ}$.
We suggested (Section 3) that the limited random distribution of three-dimensional angles $\delta$ might result from two sources of angular momentum to the accretion disks that launch the two or more pairs of jets that shaped each multipolar PN. One source has a fixed angular momentum axis, while the other adds an angular momentum with a random direction. The two sources supply angular momenta of comparable magnitude.
From the 40 PNe in our analysis, only M 2-46 is in the catalog of PNe with central binary systems built and maintained by David Jones (Jones & Boffin, 2017; Boffin & Jones, 2019; Jones, 2024). M 2-46 is different from most other multipolar PNe in that one pair is much smaller than the other and inside it (Figure 1).
Soker (2023b) examined the morphologies of the brightest PNe in the Galaxy and concluded that most, but not all, bright PNe tend to be multipolar and possess a minor to medium degree of departure from pure point symmetry. He further argued that violent binary interaction shaped the brightest PNe; triple-star interaction is not common as it leads to messy PNe, which most bright PNe do not.
We end this study by speculating on a scenario of violent binary interaction that includes a component of angular momentum with a fixed axis and one with a random axis, both of which have similar magnitudes.
Consider an AGB (or an RGB) PN progenitor with vigorous envelope convection and pulsation. We propose that the binary companion that launches the jets has an eccentric orbit and it experiences two to a few periastron passages before it enters a common envelope evolution. We assume that during and following periastron passages, the companion launches the energetic pairs of jets that shape the pairs of lobes/bubbles/ears/clumps we used to define multipolar PNe in this study. Near periastron passages, the companion accretes mass from the closest envelope parts to its orbit. The vigorous convection implies large convective cells protrude the envelope in a stochastic manner. The closest envelope part to the jet-launching companion might not be in the equatorial plane but an uprising convective cell off the equatorial plane.
Consider a protrusion at a distance $h_{c}$ from the equatorial plane. A typical value is about half the scale height $h_{c}\approx 0.5l_{P}\simeq 0.15R_{\rm G}$, whereas we take $l_{P}\simeq 0.3R_{\rm G}$ for the scale height of AGB stars with radius $R_{\rm G}$. The random component due to the convective cells has two terms. The first is due to the orbital motion and the distance of the convective protrusion from the equatorial plane (the orbit) and has a radial (negative or positive) direction
$j_{r}^{\rm conv}=h_{c}v_{\rm orb}\approx 0.15l_{P}v_{\rm orb}.$ (1)
The second term is due to the radial velocity of the convective cell protrusion, $v_{\rm conv}$; the angular momentum direction is in the negative or positive motion of the companion at periastron passage. The convective speed is the order of the sound speed in the outer envelope zones, which is somewhat below the orbital velocity at the stellar surface. The magnitude of this term is therefore
$j_{\theta}^{\rm conv}=h_{c}v_{\rm conv}\approx 0.15l_{P}v_{\rm conv}\approx 0.1l_{P}v_{\rm orb}.$ (2)
The fixed direction component is due to the orbital motion. Consider the companion moving at a distance of $\Delta a$ from the surface in the equatorial plane. It might pass even inside the outskirts of the envelope during maximum pulsational expansion or graze the envelope in the grazing envelope evolution. If the companion is inside or grazes the envelope, the density gradient supplies the net accreted angular momentum; it has a scale height of $\simeq l_{P}$, and since mass accretion occurs from both sides, we should crudely take $\Delta a\approx 0.5l_{P}$. The specific angular momentum of the fixed-axis component one
$j_{z}^{\rm orb}=\Delta a~{}v_{\rm orb}\approx 0.5l_{p}~{}v_{\rm orb},$ (3)
where we take the $z$ axis along the orbital angular momentum direction. The crude estimates of this simple scenario give the stochastic variations, equations (1) and (2), to be several tens of percent of the constant component of the specific angular momentum (equation 3); this might account for our findings of the best angle distribution to fit the observed cumulative distribution function. Future studies should explore the properties of the scenario we propose here and examine other scenarios, like accretion at the CEE’s termination or even the star’s tidal destruction at the core.
Acknowledgements
We thank Gunter Cibis and Martin Guerrero for their valuable comments. We heavily used the Planetary Nebula Image Catalogue (PNIC) of Bruce Balick (Balick, 2006) https://faculty.washington.edu/balick/PNIC/. A grant from the Pazy Foundation supported this research.